Solve the following differential equations

2) v = 2x² + y2 Ans.: y = xv4lnx+C. ху
Answer #1

2 72 х - Xyz xxx Y let 1 / 2 = v - Y=x. v Y = 8 + x duen @ by ear and ② of x cu o = 2 + 0 = x du = 3 2 odo= a dx Jode = /2